# Representative Action Selection for Large Action-Space Meta-Bandits

### v1

@@ 2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon . It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
@@
1:Input: Action space $\mathcal{A}_{full}$, Sample size $K$, Distribution of bandits.
@@
6:     Find the optimal action $a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$.
7:     $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetitions of actions are allowed
@@
We define the optimal action in a bandit instance $\theta$ as $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{full}}\mu_{a}(\theta)$.
@@
The optimal action $a^{*}(\theta)$ is unique with probability $1$ for all bandit instances.
This assumption does not exclude the case where the action space is a polytope, and there may be specific bandit instances for which the optimal actions lie on the edges, resulting in an infinite number of optimal actions. Since the bandit instances follow a continuous distribution, the probability of encountering these specific bandits is zero.
Then, define the measure $q$:
@@
The compactness assumption ensures that $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is finite and that the unique optimal action exists. Without loss of generality, let $\mathcal{A}_{full}$ be the support of measure $q$.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to the probability measure $q$. If a cluster has a higher measure under $q$, its elements are more likely to be included in the output of Algorithm 1. In fact, with high probability, Algorithm 1 outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{full}$ with respect to measure $q$. (This is a simplified version of the $\epsilon$-net theorem; see Theorem 10.2.4 of (matousek2013lectures, ).)
Lemma 3 (Measure-theoretic $\epsilon$-net).
@@ 3 Regret Analysis
For any cluster $r\in\mathcal{R}$, we define its diameter as $\operatorname{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$. Throughout the following, we set $\epsilon$ to be the maximum diameter over all clusters, i.e., $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
@@
We impose the following assumption on the partition, used only for lower bounds. This means that when the optimal action falls within an arbitrary cluster $r$, any action in this cluster is better than any action outside of it. It ensures that the clusters are relatively well-separated.
@@
For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{full}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
@@
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 4). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded using concentration property of Gaussian process. $\square$
Theorem 5 reveals that the expected regret grows linearly with the cluster diameter but logarithmically with the number of clusters. If the action space has a large overall diameter but exhibits a clustered structure—that is, actions are naturally grouped into a few tight clusters—then the reference subset can achieve significantly lower regret. In contrast, if the action space has a scattered or uniform structure, using a moderate number of clusters would lead to large cluster diameters, which in turn increases the regret of the reference subset.
As shown in the next section, Theorem 5 paves the way for analyzing Algorithm 1 by bounding its expected regret in terms of a reference subset, plus an additional term—since the proof relies on the possibility that its output matches a reference subset. Thus, the structure-regret relationship applies to Algorithm 1’s output.
@@ 3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that the algorithm may omit some clusters, whereas the reference subset includes a representative from each cluster. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to omit clusters that contribute minimally to the expected regret.
Theorem 6 (Upper bound).
@@
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
  • If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is still bounded by $\max_{a\in\mathcal{A}_{full}}\mu_{a}$. This is justified by the assumption of $0\in\mathrm{conv}(\mathcal{A})$, since $\mathbb{E}_{\theta}[\max_{a\in\mathcal{A}}\mu_{a}]$ is invariant under affine unitary transformations of $\mathcal{A}$. $\square$
Comparing Theorem 6 with Theorem 5, we observe that the algorithm-dependent upper bound exceeds that of the reference subset by an additional term:
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu^{2}_{a}\right]\right)^{1/2}.$
@@
Note that the term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ behaves similarly to entropy. To illustrate, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays rapidly when $q(r)$ is large, making the overall term negligible when $q$ is concentrated on few clusters.
In general, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{full},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{full}$:
$N(\mathcal{A}_{full},\epsilon)=\min\left\{m\in\mathbb{N}:\exists\{a_{1},\dots,a_{m}\}\subseteq\mathcal{A}_{full},\ \forall a\in\mathcal{A}_{full},\ \exists\ell,\ \|a-a_{\ell}\|_{2}\leq\epsilon\right\}.$
Theorem 7 (Partition-independent upper bound).
Under Assumption 2, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{full}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 5, and another absolute constant $c>0$,
@@
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq 2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{full},\epsilon)},\quad\text{where }K\geq c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{full},\epsilon).$
@@
The partition-independent upper bound in Theorem 7 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space. Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number—a geometric property of the action space. The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number. As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
Further, we give a corresponding lower bound of Algorithm 1:
Theorem 8 (Lower bound).
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where each cluster contains more than one action and satisfies Assumption 4. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C$ in Theorem 5 and another constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).\end{split}$
4 Examples
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{full}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{full}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{full}.$ (4)
@@ 4 Examples
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{full}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. As shown in (williams2006gaussian, ; wainwright2019high, ), the function $f$ can also be written as
@@
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (2). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{full}$.
Limitations: Gaussian assumptions may limit applicability in practice. We plan to extend this work to general random process with sub-Gaussian property.

### v2

@@ 2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by Haussler and Welzl (1986). It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
@@
1:Input: Action space $\mathcal{A}_{\operatorname*{full}}$, Sample size $K$.
@@
6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetition of actions is allowed
@@
We define the optimal action in a bandit instance $\theta$ as
$a^{*}(\theta):=\operatorname*{argmax}_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$
@@
The optimal action $a^{*}(\theta)$ is unique with probability 111 for all bandit instances.
Define the Importance Measure $q$ over a partition $\mathcal{R}$:
$\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
@@
The compactness assumption ensures that the term $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\right]$ is finite and guarantees the attainment of a unique optimal action. Without loss of generality, we assume that $\mathcal{A}_{\operatorname*{full}}$ is the support of the measure $q$.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to measure $q$. If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output. In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$. (This is a simplified version of Theorem 10.2.4 of Matousek (2013).)
Lemma 4.
@@
The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$. Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Hence, we have the following corollary:
Corollary 5.
Under the same conditions as Lemma 4, let $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, the following holds:
$\forall a\in\mathcal{A}_{\operatorname*{full}},\ \exists a^{\prime}\in\mathcal{A}\text{ such that }\|a-a^{\prime}\|_{2}<\max_{r\in\mathcal{R}}\operatorname{diam}(r).$
@@ 3 Regret Analysis
We impose the following assumption on the partition, used only for lower bounds: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it. This ensures clusters are well-separated.
@@
For any $r\in\mathcal{R}$, whenever the optimal action lies in cluster $r$, i.e., $a^{*}(\theta)\in r$, then $\mu_{a}\geq\max_{a^{\prime}\in\mathcal{A}_{\operatorname*{full}}\setminus r}\mu_{a^{\prime}},\;\forall a\in r$.
@@
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 6). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via concentration property of Gaussian process. $\square$
@@ 3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
@@
Upper Bound.
Theorem 8.
@@
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
@@
  • $r_{\ell}\cap\mathcal{A}\neq\emptyset$: The regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • $r_{\ell}\cap\mathcal{A}=\emptyset$: The regret is bounded by $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\!\mu_{a}$. $\square$
@@
The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 8 behaves similarly to entropy: it is maximized when $q$ is uniform, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Lemma 18 in Supplementary. Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
The connection between Theorem 7 and 8 is insightful:
Remark 9.
Consider placing a grid over the action space and defining a partition $\mathcal{R}$ by assigning each point in the space to its nearest grid point. In this way, the grid acts as a reference subset with respect to the partition $\mathcal{R}$. The regret upper bound in Theorem 7 applies to this grid, as it holds for any reference subset. Meanwhile, Theorem 8 applies to any partition, including $\mathcal{R}$. As a result, we obtain a regret bound for Algorithm 1 that exceeds the grid’s bound by only one additional term:
@@
$\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
where the part $\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]$ is simply a constant, as the action space $\mathcal{A}_{\operatorname*{full}}$ is fixed.
If the distribution $q$ is highly concentrated, then the expectation $\mathbb{E}_{q}[(1-q(r))^{2K}]$ is small, making the extra term negligible. In this case, the benefit of constructing an explicit grid—often a non-trivial task when the action feature vectors are unknown or high-dimensional—is limited.
On the other hand, if the measure $q$ is close to uniform, then $\min_{r\in\mathcal{R}}q(r)$ is large. In this case, with relatively high probability, Algorithm 1 outputs a geometric $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$-net (see Corollary 5). Moreover, since the partition is constructed using a grid, the cluster diameters are similar, so $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$ closely matches the spacing of the original grid—making the output resemble a slightly coarser grid.
Partition-Independent Upper Bound.
Generally speaking, the partition $\mathcal{R}$ may be unknown, and while the corresponding $q$ measure exists, it remains unspecified. We therefore provide a worst-case bound using the covering number $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, which is the smallest number of points needed to form a geometric $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big{\{}m\!\in\!\mathbb{N}:$$\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},$
$\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big{\}}.$
Theorem 10.
Under Assumption 3, there exists a point $a_{0}$ and a constant $M>0$ such that $\mathcal{A}_{\operatorname*{full}}\subset B(a_{0},M)$, a closed Euclidean ball of radius $M$ centered at $a_{0}$. Let the action space have dimension $n$, and fix a constant $0<\epsilon<M$. Let $\mathcal{A}$ be the output of Algorithm 1. For the same constant $C>0$ in Theorem 7, and another absolute constant $c>0$,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
$\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
@@
[…]
Lower Bound.
Theorem 11.
Under the same condition of Theorem 8. Let each cluster contains more than one action and satisfies Assumption 6. For the same constant $C$ in Theorem 7 and another absolute constant $c>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq&\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right)\\ &\times c\cdot\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\right)^{1/2}.\end{split}$
3.2 Regret Bound for Sub-Gaussian Process
In fact, the Gaussian process assumption can be relaxed to a sub-Gaussian one, where the random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ satisfies the following increment condition:
$\forall u>0,\quad\Pr\left[|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|_{2}^{2}}\right).$
Let $\gamma_{2}(\mathcal{S})$ denote Talagrand’s chaining functional (Talagrand, 2014) for a set $\mathcal{S}\subset\mathbb{R}^{n}$, $n\in\mathbb{N}$, equipped with the Euclidean norm. It is the sharpest known bound that for a sub-Gaussian process indexed by $\mathcal{S}$, there exists a constant $c>0$ such that $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq c\gamma_{2}(\mathcal{S})$. Using this functional, we establish a regret bound for our algorithm:
Theorem 12.
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Let $\mathcal{A}$ be output of Algorithm 1. Then, for a constant $C>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
4 Generalization and Empirical Validation
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore adopt a dimension-free view by treating the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set $\mathcal{A}_{\operatorname*{full}}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$:
$f(a):=\mu_{a},\quad\forall a\in\mathcal{A}_{\operatorname*{full}}.$
@@ 4 Generalization and Empirical Validation
where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome $f$ be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
@@
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. If the coefficients $\{\mathbf{f}_{i}\}_{i=1}^{\infty}$ are i.i.d. Gaussian, this formulation can be represented as the canonical Gaussian process model in Equation (3). Note that in this case, the actual action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.